# Engineering Carbon Credits Towards A Responsible FinTech Era: The Practices, Implications, and Future

### v1

@@ Abstract
Carbon emissions are a key factor to serious climate change, and carbon credits have emerged as an effective tool for mitigating climate deterioration and environmental damage while assisting organisations in managing their carbon footprint. In recent years, carbon credits have garnered significant attention across various economic sectors due to their substantial impact. However, fully utilizing them to address these issues remains a challenge. This study aims to enhance understanding of the engineering practices for carbon credits to develop responsible fintech solutions, thereby providing stronger incentives and insights for carbon emission management. To achieve this goal, we first conduct a comprehensive review of the negative impacts associated with organisations’ strategy of evading carbon management through non-disclosure of carbon emissions. Our findings reveal that both non-disclosure of carbon emissions and high carbon emissions negatively affect a organisation’s financial stability and market value. Therefore, it is suggested that organisations should actively manage their carbon emissions and transparently share relevant information to mitigate potential risks. This study highlights the importance of carbon credits for businesses, emphasizing their role in improving corporate sustainability. With the insights from the carbon emission, we then investigate the engineering methods that assist organisations in managing carbon emissions more cost-effectively, in particular for the data-driven computing solutions. Specifically, we analyse the factors influencing carbon prices and review the state-of-the-art carbon price prediction algorithms. These tools help organisations optimize their carbon credit purchasing strategies, thereby reducing costs and improving operational efficiency. Furthermore, we review the most recent corporate carbon emission prediction algorithms, which not only provide more accurate environmental performance assessments for investors and governments when carbon emissions data is not disclosed, but also help organisations estimate their future carbon credits needs, supporting better budget planning and strategy optimization. Finally, by integrating carbon price and carbon emission predictions, this paper proposes future research directions, including the prediction of corporate-level carbon management costs. We anticipate this will lay a foundation for future quantitative research on how carbon management practices impact corporate market value and financial performance. To the best of our knowledge, this is the first systematic review covering the various aspects of carbon credits, with a particular focus on computing solutions and engineering practices.
@@ 1 Introduction
In today’s global landscape, the Industrial Revolution mark a turning point where fossil fuels become the primary energy source, driving economic growth but also significantly increasing carbon emissions [1]. This rise in carbon emissions has led to serious climate issues, as excessive carbon dioxide forms a layer around the Earth’s surface. This layer traps more heat while reducing its dissipation, directly leading to the global temperature increases. The resulting climate change is profound and often irreversible effects on ecosystems, including glacier melting and biodiversity loss. Additionally, rising sea levels also pose a grave threat to human settlements, particularly in coastal regions. In response, the global community has invested substantial resources in promoting climate-friendly solutions [2]. A key initiative, the Paris Agreement, reflects the collective commitment of nations to limit global temperature increases to within 2 degrees Celsius by the end of this century [3].
@@
The Carbon Emissions Trading (CET) system, proposed under the Kyoto Protocol, is a key strategy to combat climate change. Introduced in 1997, CET aims to reduce greenhouse gas emissions by implementing a ‘quota trading’ mechanism. Each unit of carbon credit allows the emission of one ton of carbon dioxide [4]. Specific emission quotas are allocated to countries based on their respective development status and goals, which is then distributed to domestic organizations and organisations. Entities that exceed their allocated emissions must purchase carbon credits on the carbon trading market, while those with surplus credits can sell them on the market. This scheme has been adopted in many regions, including Europe, Asia, and the Americas [5]. In addition to regulating global carbon emissions, CET incentivises innovation in green technologies and promotes the use of renewable energy sources to further reduce carbon emissions.
@@
As the carbon credit mechanism matures, it role in mitigating global temperature rise is increasingly recognised. Decision-making must now integrate strategies for managing both carbon emissions and the carbon credits cost. For organisations with high emissions, the cost of purchasing additional carbon credits can be substantial [6], making accurate carbon credit price prediction essential for effective cost management [7].
@@
Additionally, many organisations opt to either inaccurately disclose their carbon emissions or avoid voluntarily disclosing their actual carbon emissions in practise [8]. This creates significant risks for businesses, particularly when managing carbon emission responsibilities that are closely tied to corporate reputation and financial performance. For example, some organisations fail to actively manage their carbon emissions, particularly by not purchasing carbon credits to offset excess emissions. Such practices not only clearly violate policy regulations but also expose organisations to greater compliance risks. To avoid the potential damage to their corporate image, such as fines, a decrease in market value, and harm to their reputation, many organisations choose not to disclose their actual carbon emissions, carbon management practices, or the associated operational costs in their financial reports [9]. In this paper, our findings reveal that better engineering the carbon credits for industry sectors can enable an enhanced transparency of the process, which can significantly boost the organisation performance.
@@
There are two key reasons for the increased transparency in carbon emissions disclosure. First, in the financial markets, increasing attention is being paid to whether organisations adhere to ESG (Environmental, Social, and Governance) principles. Investors, particularly institutional investors, are placing greater emphasis on organisations’ performance in sustainable development [10]. They are more inclined to allocate funds to organisations that can create long-term social value and mitigate negative externalities. This trend has significantly heightened the demand for ESG information disclosure, with investors favoring organisations that actively implement ESG principles. Consequently, financial investors typically rely on the information disclosed in organisations’ financial reports to assess their ESG performance. However, some organisations have not fully disclosed their carbon emissions and carbon management practices, making it difficult for investors to gain a comprehensive understanding of the company’s actual operations and true value. This information asymmetry not only increases uncertainty and risk in investment decisions but also potentially leads investors to miss identifying high-risk organisations. Therefore, having accurate and comprehensive information on a company’s carbon management is crucial for investors to make informed decisions.
@@
Second, government has implemented stricter regulations on organisations’ disclosure of carbon emissions. Governments also need to understand a company’s carbon emissions and management practices to formulate more effective carbon management policies and enforce penalties on non-compliant organisations, thereby supporting the achievement of emission reduction targets and controlling global temperature rise. By obtaining accurate carbon emission data, governments can better evaluate the effectiveness of current policies and adjust them in a timely manner to ensure that carbon emissions align with expected targets [11]. Additionally, understanding a company’s carbon management strategies helps governments identify industries or organisations that are lagging in emission reduction efforts, enabling the formulation of targeted support or punitive measures to promote the overall industry’s green transformation. Furthermore, governments can use this data in international climate negotiations to showcase the country’s efforts and achievements in emission reductions, thereby strengthening international confidence in the country’s climate commitments. To address the challenges faced by organisations, financial investors, and governments as outlined above, we investigate following three areas:
@@
    An analysis of the latest carbon credit price prediction algorithms and factors influencing carbon prices.
@@
    An analysis of the latest corporate carbon emission prediction algorithms.
@@
Although the existing literature increasingly focuses on the application of carbon credits, there lacks comprehensive reviews on this topic. Existing reviews tend to be limited on a specific area [2, 12, 13, 14], such as the impact of carbon credits on farmers’ income in agricultural, or the role of blockchain for carbon credits in the construction industry. However, these studies do not provide a systematic review of carbon credits from a financial and fintech perspective, leaving the broader implication for enterprises, governments, and the financial industry largely unexplored. In particular, there is a significant gap in understanding and leveraging carbon credits in fintech with advanced computing solutions.
@@
Thus, in this work, we conduct a systematic literature review [15], to comprehensively analyse the application of carbon credits in fintech sector, providing valuable insights for stakeholders in tackling climate change challenges with state-of-the-art computing algorithms and solutions. The main contributions of this article are as follows:
@@ 2 Research methodology
This study adopts a systematic literature review approach [16] to comprehensively identify and analyse the key applications of carbon credits in the fin-tech field, as well as the research progress in related areas. To ensure the transparency and replicability of this work, we strictly follow the steps, including initial literature retrieval, setting up inclusion and exclusion criteria, performing study selection process and finally conducting quality assessment.
To begin with, we consider below the key research questions for investigation:
@@
2.1 Step 1: File Retrieval
In this first step, we identified the key data sources and search terms used in the study. For the data sources, they are generally from Google scholar, Scopus, IEEE Xplore, and Science Direct. The results are collected in August 2023 and include articles from 2005 to 2024. This timeframe is chosen because key terms such as ‘carbon credits’ and ‘carbon emissions’ emerged around 2005. The search strings were constructed using key phrases from our research questions, and the search terms were combined using ‘AND’ and ‘OR’ logical operators to ensure a comprehensive search.
The selected databases are among the most widely used academic sources globally, renowned for their comprehensive and authoritative coverage. We reviewed past research papers to understand the current status and gaps in carbon credit systems, focusing on factors influencing carbon prices, algorithms for predicting carbon credit prices, algorithms for forecasting enterprise emissions, and the impact of carbon emissions disclosure on organisations.
2.2 Step 2: Inclusion and Exclusion Criteria
Herein we provide our inclusion and exclusion criteria for literature. For the inclusion criteria:
  1. 1.
    Study Types: Original research articles, conference papers, and case studies that provide new insights or data analysis to enhance understanding of the carbon credit market, price dynamics, or influencing factors.
  2. 2.
    Publication Time Frame: Considering the development of the carbon market and related policies, only literature from 2005 to present is considered to ensure relevance and timeliness of the research.
  3. 3.
    Research Focus: Literature must focus on carbon credits, carbon pricing strategies, market dynamics analysis, policy impact assessments, or the application of carbon credits in fintech.
  4. 4.
    Language: Only English-language literature is considered.
Exclusion Criteria:
  1. 1.
    Non-Peer-Reviewed Literature: Exclude literature that has not undergone peer review, such as news articles, blog posts, and informal white papers, to ensure the scientific validity and authority of the research.
  2. 2.
    Case Study Limitations: While case studies are included, those based solely on a single case and fail to provide generalised insights or in-depth analysis will be excluded.
  3. 3.
    Geographic Scope: Exclude studies focused on specific regions with known extreme immaturity of the carbon credit market or inconsistent with global trends in carbon market development.
  4. 4.
    Outdated Literature: While studies since 2005 are considered, those based on outdated or no longer applicable data, models, or assumptions are excluded.
  5. 5.
    Professional Books and Chapters: Exclude entire books and book chapters, as these are usually not peer-reviewed and are less accessible and timely than journal articles and conference papers.
  6. 6.
    Review Articles and Commentaries: Exclude review articles and commentaries that do not provide original research or new data, unless they offer unique insights into specific issues or integrate the latest developments from a large body of research.
  7. 7.
    Duplicate Studies: Exclude studies that are duplicates or have been minimally modified for different publications, to ensure the uniqueness and quality of the analysis.
By setting these detailed inclusion and exclusion criteria, our literature review section can more accurately filter the highest quality literature, which will be most relevant to our research objectives. This helps enhance the depth and breadth of the study, while also ensuring the novelty and applicability of the research findings.
2.3 Step 3: Selection Process
In the process of selecting literature, we first comprehensively considered how the literature covers various factors affecting carbon credit prices, including market demand, policy changes, economic impacts, and technological advancements. Next, we evaluated various algorithms proposed in the literature, especially those applied to predicting carbon credit prices and optimizing carbon emission trading strategies, such as machine learning, artificial intelligence, and other statistical models. We also paid special attention to the performance evaluation, test results, and practical application cases of these algorithms. Furthermore, we explored research on how to predict corporate carbon emissions and its applications, especially those discussing the accuracy and feasibility of prediction models and their applications in corporate strategic planning, risk management, and compliance reporting. Finally, we focused on literature analyzing the impact of disclosing carbon credits on corporate financial performance, brand image, market competitiveness, and sustainable development strategies.
To ensure that the selected literature meets the standards of research quality, we developed the following Quality Assessment (QA) criteria:
  • QA1: Does the literature qualitatively and quantitatively explore the impact of non-disclosure of carbon emissions and carbon management information?
  • QA2: Does the literature conduct an in-depth analysis of factors influencing carbon credit prices, providing sufficient evidence and discussion?
  • QA3: Does the literature propose new methods or algorithms for predicting carbon credit prices and validate their effectiveness?
  • QA4: Does the literature propose new methods or algorithms for predicting corporate carbon emission and validate their effectiveness?
  • QA5: Does the literature clearly describe the research results and provide a thorough evaluation of models or algorithms?
In summary, we have collected a comprehensive set of research articles focusing on the different aspects of carbon credit systems, including factors influencing carbon prices, prediction algorithms for carbon credit prices and corporate emissions, and the impact of carbon emissions disclosure on organisations. This systematic approach ensures that our study incorporates high-quality literature, providing valuable insights into the dynamics of the carbon credit market and its intersection with financial technologies.
3 For FinTech: The Impact of Disclosing Carbon Emissions on organisations
@@
Studying carbon emission disclosure practices in these countries is of significant importance. First, differences in policies and market environments regarding carbon emission management and disclosure across countries can reveal the factors that drive or hinder organisations from disclosing carbon emission information. Second, an in-depth exploration of the impact of carbon emission disclosure can help understand how organisations can enhance market trust through transparent environmental responsibility practices and provide valuable insights for policymakers to further improve carbon credit trading mechanisms.
@@
Recent studies in the United States suggest that firms with voluntary disclosure of carbon emissions experience higher valuations than their counterparts without voluntary disclosure. For example, a study conducted in the United States concluded that the market imposes stricter penalties on those who choose not to disclose relevant information, and that high carbon emissions have a negative impact on a company’s market value. [17].
@@
Similar results have been replicated in the United Kingdom. A research in the United Kingdom, another developed country, shows that carbon emissions significantly negatively reduce corporate value, with markets typically reacting unfavorably to organisations with high carbon emissions. Even when organisations disclose more carbon emission information, the market response remains negative. This finding aligns with the results observed in the U.S. market [10]. Furthermore, the study have revealed a significant positive correlation between capital expenditure and carbon emission disclosure. Specifically, the more capital expenditure a company has, the more carbon emission information it discloses in its annual reports. This relationship is likely due to the fact that capital expenditure is typically associated with increased production activities, which in turn lead to higher carbon emissions [10].
@@
Additionally, the study highlights that robust internal corporate governance, such as board size, board independence, and gender diversity enhances the level of carbon emission disclosure [10]. Other studies focusing on UK organisations have explored the relationship between carbon emission, disclosure, and financial performance. The findings indicate a negative correlation between carbon emission levels and corporate financial performance. Conversely, a significant positive correlation exists between carbon information disclosure and financial performance, indicating that organisations that voluntarily disclose carbon information experience significant improvements in financial performance [18, 19].
@@
In the Australian market, disclosure policy plays a key role in influencing voluntary disclosure. The introduction of the National Greenhouse and Energy Reporting (NGER) Act in 2007 boosted voluntary disclosures of carbon emission, especially among larger, more visible firms. The research found that company size and carbon emission levels are key determinants of the extent of carbon emission disclosure. Larger organisations tend to disclose more carbon emission information, because they are more visible to the public and the regulators, thus having greater motivation to meet societal expectations through transparent disclosure practices. Furthermore, organisations with higher carbon emissions are more inclined to provide detailed disclosures, particularly in high-emission industries such as energy, transportation, materials, and utilities, where organisations face greater social and regulatory pressures, leading to more proactive and detailed disclosures [8].
@@
The study also suggests that good corporate governance is crucial in promoting carbon emission disclosure. organisations with better governance structures (such as strong board independence and high governance scores) are more likely to disclose comprehensive carbon emission information, reflecting the positive role of good governance in enhancing corporate environmental transparency. Moreover, significant differences in disclosure practices exist across industries. organisations in high-emission industries, due to greater social and political pressures, generally disclose more detailed carbon emission information, while in low-emission industries, better disclosure practices are often observed in larger organisations or those with more robust governance structures [8].
@@
Research in Japan further supports the aforementioned conclusions. Data analysis shows that reducing carbon emissions and disclosing carbon management activities have a positive impact on corporate market value. Specifically, the study reveals a significant negative correlation between corporate carbon emission levels and stock market value, while carbon management disclosure is positively correlated with stock market value. Additionally, the study finds that the positive correlation between carbon management disclosure and stock market value becomes more pronounced when carbon emission levels are higher. This suggests that reducing carbon emissions and actively disclosing related management activities are critical to a company’s sustainable development [20].
@@
In contrast to developed countries, studies conducted in developing countries have different results. For instance, a study in Indonesia found no direct effect of carbon emissions disclosure on firm value. Investors in these markets tend to prioritize financial performance over environmental performance. However, the study also found that disclosing carbon emission information can improve corporate financial performance, as organisations that disclose such information can attract environmentally conscious consumers, thereby increasing revenue and profitability, which is consistent with observations in developed countries [11].
@@
Further analysis suggests that effective government policy guidance is crucial for addressing climate change and enabling capital markets to hold organisations accountable for their carbon emissions. The implementation of mandatory disclosure policies, requiring organisations to publicly disclose their carbon emissions, is a key step in this process. Under such regulatory environments, the negative impact of carbon emissions on corporate market value becomes more significant, while voluntary disclosure of carbon emission information can partially mitigate this impact. It is through such mechanisms that the market can effectively penalize organisations with high carbon emissions and exert further pressure on those that choose not to disclose relevant information, which is why developed country governments generally adopt mandatory disclosure policies [11].
@@
For example, a study in Malaysia found that corporate carbon emission levels and disclosure do not significantly impact market value, further indicating that in the absence of mandatory policies, organisations lack motivation to strengthen carbon management [21]. Similar findings were observed in South Korea, where the impact of carbon emissions and disclosure on corporate market value is limited, as the capital market in Korea places greater emphasis on corporate profitability rather than environmental impact. Therefore, effective government policy guidance is particularly important, requiring stricter regulations to direct market attention towards corporate carbon emission behaviors [22]. In contrast, a study in the UK demonstrated that the government’s mandatory disclosure requirements led to greater negative impacts on the market value of high-carbon-emission organisations, further confirming the critical role of policy in driving corporate carbon management and information disclosure [23].

### v2

# These $\equalcontThese authors contributed equally to this work.$

@@ Abstract
Carbon emissions are a key factor to serious climate change, and carbon credits have emerged as an effective tool for mitigating climate deterioration and environmental damage while assisting organisations in managing their carbon footprint. In recent years, carbon credits have garnered significant attention across various economic sectors due to their substantial impact. However, fully utilizing them to address these issues remains a challenge. This study aims to enhance understanding of the engineering practices for carbon credits to develop responsible fintech solutions, thereby providing stronger incentives and insights for carbon emission management. To achieve this goal, we first conduct a comprehensive review of the negative impacts associated with organisations’ strategy of evading carbon management through non-disclosure of carbon emissions. Our findings reveal that both non-disclosure of carbon emissions and high carbon emissions negatively affect a organisation’s financial stability and market value. Therefore, it is suggested that organisations should actively manage their carbon emissions and transparently share relevant information to mitigate potential risks. This study highlights the importance of carbon credits for businesses, emphasizing their role in improving corporate sustainability. With the insights from the carbon emission, we then investigate the engineering methods that assist organisations in managing carbon emissions more cost-effectively, in particular for the data-driven computing solutions. Specifically, we analyse the factors influencing carbon prices and review the state-of-the-art carbon price prediction algorithms. These tools help organisations optimize their carbon credit purchasing strategies, thereby reducing costs and improving operational efficiency. Furthermore, we review the most recent corporate carbon emission prediction algorithms, which not only provide more accurate environmental performance assessments for investors and governments when carbon emissions data is not disclosed, but also help organisations estimate their future carbon credits needs, supporting better budget planning and strategy optimization. Finally, by integrating carbon price and carbon emission predictions, this paper proposes future research directions, including the prediction of corporate-level carbon management costs. We anticipate this will lay a foundation for future quantitative research on how carbon management practices impact corporate market value and financial performance. To the best of our knowledge, this is the first systematic review covering the various aspects of carbon credits, with a particular focus on computing solutions and engineering practices. By enhancing transparency through advanced computing solutions, this research underlines the role of AI in fostering social accountability, ensuring that the transition to a low-carbon economy remains inclusive and trustworthy.
keywords:
Carbon price, carbon emission, FinTech, machine learning, artificial intelligence
@@ 1 Introduction
In today’s global landscape, the Industrial Revolution mark a turning point where fossil fuels become the primary energy source, driving economic growth but also significantly increasing carbon emissions [carbon1]. This rise in carbon emissions has led to serious climate issues, as excessive carbon dioxide forms a layer around the Earth’s surface. This layer traps more heat while reducing its dissipation, directly leading to the global temperature increases. The resulting climate change is profound and often irreversible effects on ecosystems, including glacier melting and biodiversity loss. Additionally, rising sea levels also pose a grave threat to human settlements, particularly in coastal regions. These environmental shifts are not merely technical or economic challenges, but profound social crises that necessitate a re-evaluation of the ethical contract between technology, finance, and society. In response, the global community has invested substantial resources in promoting climate-friendly solutions [carbon2]. A key initiative, the Paris Agreement, reflects the collective commitment of nations to limit global temperature increases to within 2 degrees Celsius by the end of this century [carbon3].
@@
The Carbon Emissions Trading (CET) system, proposed under the Kyoto Protocol, is a key strategy to combat climate change. Introduced in 1997, CET aims to reduce greenhouse gas emissions by implementing a ‘quota trading’ mechanism. Each unit of carbon credit allows the emission of one ton of carbon dioxide [carbon4]. Specific emission quotas are allocated to countries based on their respective development status and goals, which is then distributed to domestic organizations and organisations. Entities that exceed their allocated emissions must purchase carbon credits on the carbon trading market, while those with surplus credits can sell them on the market. This scheme has been adopted in many regions, including Europe, Asia, and the Americas [carbon5]. In addition to regulating global carbon emissions, CET incentivises innovation in green technologies and promotes the use of renewable energy sources to further reduce carbon emissions.
@@
As the carbon credit mechanism matures, it role in mitigating global temperature rise is increasingly recognised. Decision-making must now integrate strategies for managing both carbon emissions and the carbon credits cost. For organisations with high emissions, the cost of purchasing additional carbon credits can be substantial [carbonmanagement], making accurate carbon credit price prediction essential for effective cost management [carbonprice].
@@
Additionally, many organisations opt to either inaccurately disclose their carbon emissions or avoid voluntarily disclosing their actual carbon emissions in practise [carbon69]. This creates significant risks for businesses, particularly when managing carbon emission responsibilities that are closely tied to corporate reputation and financial performance. For example, some organisations fail to actively manage their carbon emissions, particularly by not purchasing carbon credits to offset excess emissions. Such practices not only clearly violate policy regulations but also expose organisations to greater compliance risks. To avoid the potential damage to their corporate image, such as fines, a decrease in market value, and harm to their reputation, many organisations choose not to disclose their actual carbon emissions, carbon management practices, or the associated operational costs in their financial reports [carbon72]. In this paper, our findings reveal that better engineering the carbon credits for industry sectors can enable an enhanced transparency of the process, which can significantly boost the organisation performance.
@@
There are two key reasons for the increased transparency in carbon emissions disclosure. First, in the financial markets, increasing attention is being paid to whether organisations adhere to ESG (Environmental, Social, and Governance) principles. Investors, particularly institutional investors, are placing greater emphasis on organisations’ performance in sustainable development [carbon68]. They are more inclined to allocate funds to organisations that can create long-term social value and mitigate negative externalities. This trend has significantly heightened the demand for ESG information disclosure, with investors favoring organisations that actively implement ESG principles. Consequently, financial investors typically rely on the information disclosed in organisations’ financial reports to assess their ESG performance. However, some organisations have not fully disclosed their carbon emissions and carbon management practices, making it difficult for investors to gain a comprehensive understanding of the company’s actual operations and true value. This information asymmetry not only increases uncertainty and risk in investment decisions but also potentially leads investors to miss identifying high-risk organisations. Therefore, having accurate and comprehensive information on a company’s carbon management is crucial for investors to make informed decisions. Furthermore, the integration of AI into carbon credit engineering transcends corporate profit-seeking; it empowers public oversight and reinforces social trust by mitigating the risks of greenwashing. This alignment of data-driven precision with collective environmental welfare is fundamental to the emergence of a truly responsible FinTech era.
@@
Second, government has implemented stricter regulations on organisations’ disclosure of carbon emissions. Governments also need to understand a company’s carbon emissions and management practices to formulate more effective carbon management policies and enforce penalties on non-compliant organisations, thereby supporting the achievement of emission reduction targets and controlling global temperature rise. By obtaining accurate carbon emission data, governments can better evaluate the effectiveness of current policies and adjust them in a timely manner to ensure that carbon emissions align with expected targets [carbon70]. Additionally, understanding a company’s carbon management strategies helps governments identify industries or organisations that are lagging in emission reduction efforts, enabling the formulation of targeted support or punitive measures to promote the overall industry’s green transformation. Furthermore, governments can use this data in international climate negotiations to showcase the country’s efforts and achievements in emission reductions, thereby strengthening international confidence in the country’s climate commitments. To address the challenges faced by organisations, financial investors, and governments as outlined above, we investigate following three areas:
@@
    The latest carbon credit price prediction algorithms and factors influencing carbon prices.
@@
    The latest corporate carbon emission prediction algorithms.
@@
Although the existing literature increasingly focuses on the application of carbon credits, there lacks comprehensive reviews on this topic. Existing reviews tend to be limited on a specific area [carbon2, carbon74, carbon75, carbon76], such as the impact of carbon credits on farmers’ income in agricultural, or the role of blockchain for carbon credits in the construction industry. However, these studies do not provide a systematic review of carbon credits from a financial and fintech perspective, leaving the broader implication for enterprises, governments, and the financial industry largely unexplored. In particular, there is a significant gap in understanding and leveraging carbon credits in fintech with advanced computing solutions.
@@
Thus, in this work, we conduct a systematic literature review [carbon10], to comprehensively analyse the application of carbon credits in fintech sector, providing valuable insights for stakeholders in tackling climate change challenges with state-of-the-art computing algorithms and solutions. The main contributions of this article are as follows:
@@ 2 Research methodology
2.1 Research questions
This study adopts a systematic literature review approach [SLR] to comprehensively identify and analyse the key applications of carbon credits in the fin-tech field, as well as the research progress in related areas. To ensure the transparency and replicability of this work, we strictly follow the steps, including initial literature retrieval, setting up inclusion and exclusion criteria, performing study selection process and finally conducting quality assessment. We consider following research questions as an outline to initiate our investigation:
@@ 2.1 Research questions
2.2 Methodology
We implement this work as a systematic literature review, with an inclusion of a broad range of data sources and defined keyword search string. We aims to cover as many impactful paper as possible by soliciting the results from the sources from Google scholar, Scopus, IEEE Xplore, and Science Direct. The results are collected till February 2025 and include articles from 2005 to 2024, since key terms such as ‘carbon credits’ and ‘carbon emissions’ emerge. The search strings were constructed using key phrases from our research questions, and the search terms were combined using ‘AND’ and ‘OR’ operators to ensure an exhaustive search. We reviewed past research papers to understand the current status and gaps in carbon credit systems, focusing on factors influencing carbon prices, algorithms for predicting carbon credit prices, algorithms for forecasting enterprise emissions, and the impact of carbon emissions disclosure on organizations.
To facilitate this process, we set up several inclusion and exclusion criteria. For the inclusion criteria, we consider peer-reviewed English-only original research articles, conference papers, and case studies that provide new insights or data analysis to enhance understanding of the carbon credit market, price dynamics, or influencing factors. For the exclusion criteria, we remove the duplicate studies to ensure the uniqueness and quality of the analysis. We also exclude studies focused on specific regions with known extreme immaturity of the carbon credit market or inconsistent with global trends in carbon market development.
In total, we retrieved a total of 59,000 relevant papers from various databases. During the initial screening phase, we reviewed the titles, abstracts and keywords of the papers, selecting 205 potentially relevant studies. Subsequently, we conducted a detailed full-text review of these remaining papers and applied our inclusion criteria for further filtering. In the end, we identified 75 papers directly related to the theme of this study. These selected papers form the foundation for the subsequent analysis and review.
3 For Finance: The Impact of Disclosing Carbon Emissions on organisations
@@
Studying carbon emission disclosure practices in these countries is of significant importance. First, differences in policies and market environments regarding carbon emission management and disclosure across countries can reveal the factors that drive or hinder organisations from disclosing carbon emission information. Second, an in-depth exploration of the impact of carbon emission disclosure help understand how organisations can enhance market trust through transparent environmental responsibility practices and provide valuable insights for policymakers to further improve carbon credit trading mechanisms.
@@
Recent studies in the United States suggest that firms with voluntary disclosure of carbon emissions experience higher valuations than their counterparts without voluntary disclosure. For example, a study conducted in the United States concluded that the market imposes stricter penalties on those who choose not to disclose relevant information, and that high carbon emissions have a negative impact on a company’s market value [carbon67].
@@
Similar results have been replicated in the United Kingdom. A research in the United Kingdom, another developed country, shows that carbon emissions significantly negatively reduce corporate value, with markets typically reacting unfavorably to organisations with high carbon emissions. Even when organisations disclose more carbon emission information, the market response remains negative. This finding aligns with the results observed in the U.S. market [carbon68]. Furthermore, the study have revealed a significant positive correlation between capital expenditure and carbon emission disclosure. Specifically, the more capital expenditure a company has, the more carbon emission information it discloses in its annual reports. This relationship is likely due to the fact that capital expenditure is typically associated with increased production activities, which in turn lead to higher carbon emissions [carbon68].
@@
Additionally, the study highlights that robust internal corporate governance, such as board size, board independence, and gender diversity enhances the level of carbon emission disclosure [carbon68]. Other studies focusing on UK organisations have explored the relationship between carbon emission, disclosure, and financial performance. The findings indicate a negative correlation between carbon emission levels and corporate financial performance. Conversely, a significant positive correlation exists between carbon information disclosure and financial performance, indicating that organisations that voluntarily disclose carbon information experience significant improvements in financial performance [carbon77, carbon81].
@@
In the Australian market, disclosure policy plays a key role in influencing voluntary disclosure. The introduction of the National Greenhouse and Energy Reporting (NGER) Act in 2007 boosted voluntary disclosures of carbon emission, especially among larger, more visible firms. The research found that company size and carbon emission levels are key determinants of the extent of carbon emission disclosure. Larger organisations tend to disclose more carbon emission information, because they are more visible to the public and the regulators, thus having greater motivation to meet societal expectations through transparent disclosure practices. Furthermore, organisations with higher carbon emissions are more inclined to provide detailed disclosures, particularly in high-emission industries such as energy, transportation, materials, and utilities, where organisations face greater social and regulatory pressures, leading to more proactive and detailed disclosures [carbon69].
@@
The study also suggests that good corporate governance is crucial in promoting carbon emission disclosure. organisations with better governance structures (such as strong board independence and high governance scores) are more likely to disclose comprehensive carbon emission information, reflecting the positive role of good governance in enhancing corporate environmental transparency. Moreover, significant differences in disclosure practices exist across industries. organisations in high-emission industries, due to greater social and political pressures, generally disclose more detailed carbon emission information, while in low-emission industries, better disclosure practices are often observed in larger organisations or those with more robust governance structures [carbon69].
@@
Research in Japan further supports the aforementioned conclusions. Data analysis shows that reducing carbon emissions and disclosing carbon management activities have a positive impact on corporate market value. Specifically, the study reveals a significant negative correlation between corporate carbon emission levels and stock market value, while carbon management disclosure is positively correlated with stock market value. Additionally, the study finds that the positive correlation between carbon management disclosure and stock market value becomes more pronounced when carbon emission levels are higher. This suggests that reducing carbon emissions and actively disclosing related management activities are critical to a company’s sustainable development [carbon73].
@@
In contrast to developed countries, studies conducted in developing countries have different results. For instance, a study in Indonesia found no direct effect of carbon emissions disclosure on firm value. Investors in these markets tend to prioritize financial performance over environmental performance. However, the study also found that disclosing carbon emission information can improve corporate financial performance, as organisations that disclose such information can attract environmentally conscious consumers, thereby increasing revenue and profitability, which is consistent with observations in developed countries [carbon70].
@@
Further analysis suggests that effective government policy guidance is crucial for addressing climate change and enabling capital markets to hold organisations accountable for their carbon emissions. The implementation of mandatory disclosure policies, requiring organisations to publicly disclose their carbon emissions, is a key step in this process. Under such regulatory environments, the negative impact of carbon emissions on corporate market value becomes more significant, while voluntary disclosure of carbon emission information can partially mitigate this impact. It is through such mechanisms that the market can effectively penalize organisations with high carbon emissions and exert further pressure on those that choose not to disclose relevant information, which is why developed country governments generally adopt mandatory disclosure policies [carbon70].
@@
For example, a study in Malaysia found that corporate carbon emission levels and disclosure do not significantly impact market value, further indicating that in the absence of mandatory policies, organisations lack motivation to strengthen carbon management [carbon71]. Similar findings were observed in South Korea, where the impact of carbon emissions and disclosure on corporate market value is limited, as the capital market in Korea places greater emphasis on corporate profitability rather than environmental impact. Therefore, effective government policy guidance is particularly important, requiring stricter regulations to direct market attention towards corporate carbon emission behaviors [carbon86]. In contrast, a study in the UK demonstrated that the government’s mandatory disclosure requirements led to greater negative impacts on the market value of high-carbon-emission organisations, further confirming the critical role of policy in driving corporate carbon management and information disclosure [carbon85].